Previous

Calendar

April

2008

Next
Sunday Monday Tuesday Wednesday Thursday Friday Saturday

30

JD90

31

JD91

1

JD92

2

JD93 (warm, calm, very wet from melting snow)
Site Visit (JB, SL, NT, Nate M, Chris): Details

downloaded from soil, tcs and noah loggers, and swapped pcmcia cards at the mrs logger

see details for snow depths swapped 4 flasks at the west robot

swapped N2 cylinders at the tower (see details)

dug out the lowest prof inlet on the mini tower

3

JD94

4

JD95

5

JD96

6

JD97

7

JD98

8

JD99

9

JD100

10

JD101
A few short power-outages at tower

11

JD102

12

JD103

13

JD104

14

JD105

15

JD106 (warm/blustery)
Site Visit (JB, SpB, JH): Details , Photos (33)

at 10:30 MST, fixed power connection for NCTC logger

at 11:35 MST, swiped krypton hygrometer

at 11:45 MST, moved 12.5 li-7500 closer to ati sonic

from ~11:50-11:55 MST, adjusted flow rates in the LI-6262 sample cell

Collected soil samples (soil still frozen).

Jeep Rental

16

JD107

17

JD108

18

JD109 (warm, breezy, wet/heavy snow)
Site Visit (JB): Details

swapped 4 flasks at west robot, hauled out suitcase to ship to Utah

hauled empty N2 cylinder (CC193045) back to trailer

see details for snow depth info

span @ 1250 psi, N2 @ 1325 psi

swapped desiccant at the LI-6251, see details

downloaded from soil logger, see details

put the SC32B in Noah's logger box, see details

dug out the lowest prof inlet

19

JD110

20

JD111

21

JD112

22

JD113
Site Visit (JH): Details

Collected soil samples---soil still frozen.

23

JD114

24

JD115

25

JD116

26

JD117

27

JD118

28

JD119
Site Visit (JH): Details

Collected soil samples---soil starting to thaw out.

29

JD120

30

JD121

1

JD122

2

JD123

3

JD124